OpenStudy (anonymous):

Does (sin^3x+cos^3x)/(sinx+cosx) = 1- sinxcosx ??

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Luigi0210

OpenStudy (loser66):

yup

OpenStudy (anonymous):

can you prove it for me ?

OpenStudy (anonymous):

Im having trouble

OpenStudy (loser66):

just apply a^3 + b^3 =(a+b)(a^2-ab+b^2) where a = sin, b = cos

OpenStudy (loser66):